OpenStudy (anonymous):

Show that d/dx [(2x)/√(x+1)] = [(x+2)/(x+1)^(3/2)]. Hence evaluate Intergration from 0 to 8 [(x+2)/(x+1)^(3/2)] dx.

hartnn (hartnn):

do you know the quotient rule in derivatives ?

OpenStudy (anonymous):

Yes

hartnn (hartnn):

so did you try it ? d/dx [(2x)/√(x+1)] = ... ?

OpenStudy (anonymous):

Nope, I wasn't sure what to do. Hold on, I'll try it and get back to you .Thanks :)

hartnn (hartnn):

sure, take your time :)

OpenStudy (anonymous):

It equals [(x+2)/(x+1)^(3/2)]

hartnn (hartnn):

so you could successfully do the first part, right ? how about 2nd part, any ideas ?

OpenStudy (anonymous):

So am I supposed to integrate the value I just found?

hartnn (hartnn):

no, integration is anti-derivative! \(\Large if, \dfrac{d}{dx}f(x) = F(x) \\\Large then, \int F(x) = f(x) +c\)

hartnn (hartnn):

so integration of [(x+2)/(x+1)^(3/2)] dx is just (2x)/√(x+1) makes sense ?

OpenStudy (anonymous):

It's just the reverse right?

hartnn (hartnn):

yup, thats correct :)

hartnn (hartnn):

and we first plug in upper limit, then, the lower limit so plug in x =8 in [(2x)/√(x+1)] then x =0

OpenStudy (anonymous):

Then proceed to integrate it?

hartnn (hartnn):

what? you need not integrate anything here

hartnn (hartnn):

just because derivative of [(2x)/√(x+1)] = [(x+2)/(x+1)^(3/2)] so integration of [(x+2)/(x+1)^(3/2)] is [(2x)/√(x+1)]

OpenStudy (anonymous):

Oh okay, I get it now. I just need to sub in the x values as the limits?

hartnn (hartnn):

yes, thats right

hartnn (hartnn):

and then subtract them

OpenStudy (anonymous):

Alright thank you :D Appreciate your help

hartnn (hartnn):

welcome ^_^